OpenStudy (anonymous):

What is the slope of this equation? y = -x + 3

OpenStudy (johnweldon1993):

slope intercept form says y = mx + b where m is the slope...and b is the y-intercept so you have y = -x + 3 y = mx + b you're slope would be?

OpenStudy (johnweldon1993):

just in case *hint .... you can always put a 1 coefficient in front of 'x' ....because 1x...is the same as x correct? so now you can have y = -1x + 3 y = mx + b remember 'm' is slope....so what is your slope?

OpenStudy (anonymous):

-1 yeah? @johnweldon1993

OpenStudy (johnweldon1993):

that is correct! :)

OpenStudy (anonymous):

Thank youuu (:

OpenStudy (johnweldon1993):

anytime :)